OpenStudy (fairegaming):

What is... z/7 / 1/7 So its z over 7 divided by 1 over 7.

OpenStudy (anonymous):

I believe it's Z

OpenStudy (anonymous):

Division of fractions can be transformed into a multiplication of fractions by multiplying the numerator by the reciprocal of the denominator. For example,\[\frac{ \frac{ x }{ 3 } }{ \frac{ 1 }{ 6 } }=\frac{ x }{ 3 }\times \frac{ 6 }{ 1 } = \frac{ 6x }{ 3 } = 2x\]

OpenStudy (anonymous):

how many one somethings are in 8 somethings? 8 how many one sevenths in five sevenths? 5 how many 1 sevenths in z sevenths?

OpenStudy (fairegaming):

So its Z???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Correct. Good job.

OpenStudy (fairegaming):

Thanks

OpenStudy (anonymous):

You're welcome